# All That Glitters is Not Novel: Plagiarism in AI Generated Research

### v1

@@ Acknowledgements
Appendix A Additional Case Study
@@
As shown in Figure 4, the proposed methodology exhibits clear similarities with the original paper. The LLM-generated paper proposes combining outputs from two diffusion models for lower and higher resolution data using learned weights, concepts previously explored in different contexts—combining multiple diffusion paths in Park et al. (2024) and jointly training diffusion models at multiple resolutions in Gu et al. (2023). Each major component of the proposed method has a corresponding match in Park et al. (2024): the gating mechanism is identical, and the “diversity loss” is closely analogous to “diffusion prior loss” from the original work, both utilizing pair-wise distance functions.
The methodology for generating these research papers involves $20$ rounds of iterative search using LLM and Semantic Scholar to identify relevant citations (Lu et al., 2024a). However, this process fails to identify and cite the original work (Park et al., 2024). This oversight demonstrates important limitations in using current methods to find closely related work in LLM-generated research documents, which in turn reveals their inadequacy for automated plagiarism detection. The case also illustrates how generated content can reformulate existing technical approaches using different terminology while maintaining the same underlying methodology.
This case, along with our main example 5.2, reveals a pattern where generated content appears to systematically reformulate existing research while maintaining core methodological similarities that are difficult to detect through conventional review processes.
@@
Appendix B Example Case from Synthetic Dataset

### v2

@@ Acknowledgements
Appendix A Case Study 2
@@
As shown in Figure 4, the proposed methodology exhibits clear similarities with the original paper. The LLM-generated paper proposes combining outputs from two diffusion models for lower and higher resolution data using learned weights, concepts previously explored—combining multiple diffusion paths in Park et al. (2024) and jointly training diffusion models at multiple resolutions in Gu et al. (2023). Each major component of the proposed method has a corresponding match in Park et al. (2024): the gating mechanism is identical, and the “diversity loss” is closely analogous to “diffusion prior loss” from the original work, both utilizing pair-wise distance functions.
The methodology for generating these research papers involves $20$ rounds of iterative search using LLM and Semantic Scholar to identify relevant citations (Lu et al., 2024a). However, this process fails to identify and cite the original work (Park et al., 2024). This oversight demonstrates limitations in using current methods to find closely related work in LLM-generated research documents, which in turn reveals their inadequacy for automated plagiarism detection. This case also illustrates how generated content can reformulate existing technical approaches using different terminology while maintaining the same underlying methodology.
@@
Appendix B Case Study 3: AI-Generated ICLR Workshop Paper
We examine an AI-generated paper titled “Compositional Regularization: Unexpected Obstacles in Enhancing Neural Network Generalization” that received scores of 6, 7, 6 at an ICLR 2025 workshop—above the average acceptance threshold. This paper was generated using the AI-Scientist-v2 system described in Yamada et al. (2025). We discover substantial similarity to an existing work that was not cited in the AI-generated paper.
The AI-generated paper’s core contribution, termed “compositional regularization,” is identical to the $(\Delta h_{t})^{2}$ regularization term that was evaluated in Table 3 of the “Regularizing RNNs by Stabilizing Activations” paper Krueger and Memisevic (2015). The original authors found this formulation less effective than their proposed norm-stabilizer approach, which aligns with the negative results reported in the AI-generated paper. Notably, the AI-generated paper provides no theoretical justification for why penalizing changes in hidden states should enhance compositional generalization.
The AI-generated paper essentially borrowed the core contribution from previous work without attribution and applied it to an unsuitable domain where it neither theoretically nor practically succeeds. We estimate the similarity score at $3$ or $4$ on our evaluation scale. Krueger and Memisevic (2015) confirmed that the AI-generated work is “definitively not novel” and very similar to one of their ablations and the concept of “slow features” Jonschkowski and Brock (2015), both of which are not cited in the AI-generated paper.
This similarity was not detected during the review process, nor by the AI-Scientist-v2 team’s own plagiarism detection measures. This case provides additional evidence supporting our paper’s central argument: AI-generated research can borrow specific ideas from prior work without proper attribution in ways that are difficult for humans and current plagiarism detection systems to identify.
Appendix C Example Case from Synthetic Dataset
@@ 3. Key Similarities and Differences:
Peer Review Analysis Prompt
Please analyze this academic review text and assign a plagiarism score from 1-5 using this rubric:
5 - Direct Copy: One-to-one mapping between methodology and existing methods in one/two closely related prior papers
4 - Combined Borrowing: Significant portion is mix-and-match from 2-3 prior works
3 - Partial Overlap: Bears decent similarity with existing methods, but no exact correspondence
2 - Minor Similarity: Very slight resemblance with existing papers. Mostly novel
1 - Original: Completely novel
Important: For scores 4 or 5, the review MUST specifically mention the prior papers that were allegedly copied from. Generic comments about missing citations or similarity without naming specific papers should not result in scores of 4 or 5.
If the review seems incomplete or nonsensical or not related or if it’s actually an author response, please assign a score of 1 to be conservative.
Review text: {review_text}
Please provide a brief explanation of your reasoning, followed by ONLY the numeric score (1-5) on a new line (no punctuation, quotes or anything else).
Table 11: The prompt used to instruct Claude 3.7 Sonnet to analyze peer reviews and assign plagiarism scores ($1$-$5$) for human-written papers based on reviewer comments.